Ellora wants to accumulate $150 000.00 in an RRSP by making annual contributions of $5000.00 at the end of each year. If interest is 5.5% compounded yearly, calculate how long she has to make contributions. Select one: a. n = 18.202125 b. n = 18.086955 c. n = 17.585794 d. n = 17.462170 e. n = 17.489212
Answer #1

Future value of annuity=Annuity[(1+rate)^time period-1]/rate

150,000=5000[(1.055)^n-1]/0.055

150,000=90909.0909[(1.055)^n-1]

[(1.055)^n-1]=(150,000/90909.0909)

(1.055)^n=(150,000/90909.0909)+1

(1.055)^n=2.65

Taking log on both sides;

n*log 1.055=log 2.65

n=log 2.65/log 1.055

=18.202125(Approx)